OpenStudy (anonymous):

Solve 4 log12 2 + log12 x = log12 96. (1 point) x = 88 x = 80 x = 12 x = 6

OpenStudy (anonymous):

@coolgirl12

Nnesha (nnesha):

12 is base right ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\huge\rm 4\log_{12} 2 + \log_{12} x = \log_{12} 96\] is this ur question ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

okay\[\large \color{green}{\rm 2 \log y = \log y^2}\] do you know this rules \[\huge\rm log ~x + \log~ y = \log ( x \times y )\]

Nnesha (nnesha):

so use that rules and change number which is multiplying by log change that to exponent of 2** left side 4 log 2 = ???

Nnesha (nnesha):

did you understand ??